OpenStudy (anonymous):

Identify a pattern and find the next number in the pattern. –2, 4, 10, 16

OpenStudy (anonymous):

22

OpenStudy (anonymous):

16- 6 = 10 10-6 = 4 4-6 = -2

OpenStudy (anonymous):

Do you understand ?

OpenStudy (anonymous):

wait let me study it

OpenStudy (anonymous):

Now we do the opposite to find what the next number will be. So instead of -6) it will be (+6) So..... 16+6= 22 :)

OpenStudy (anonymous):

why did we start off with minus, and change to adding?

OpenStudy (anonymous):

Let's start from the other end.

OpenStudy (anonymous):

-2 + (6) = 4 4 + (6) = 10 10 + (6) = 16 12 + (6) = 22

OpenStudy (anonymous):